OpenStudy (anonymous):

. In the function f(x) = 3x + 12 how does a decrease of 1 in the independent variable x affect f(x)? f(x) will have an increase of 3x + 3 f(x) will have an increase of 3 f(x) will have a decrease of 3 f(x) will have a decrease of 15

OpenStudy (anonymous):

Substitute -1 in place of x yields; f(-1)=3(-1)+12=-3+12=9 Hence an increase of 9... Which doesn't seem to be an answer...

OpenStudy (anonymous):

Are you sure you've typed in the question correctly?

OpenStudy (anonymous):

let me see